# Political Neutrality in AI is Impossible — But Here is How to Approximate it

### v1

@@ 3.2 System-Level
Tradeoffs. The greatest benefit of reflective neutrality is the enhancement of user agency and utilization, as the system is tailored to align with the user’s specific wants and needs. Additionally, it allows system developers to avoid a one-size-fits-all bias, instead curating the bias to suit the individual end-user. However, it can be argued that personalized systems may reinforce users’ inherent biases (Ludwig et al., 2023), potentially causing harm by reducing their exposure to opposing viewpoints (Pariser, 2011).
@@
System transparency can be achieved through thorough documentation of potential political bias, its sources, and manifestations. More specifically, AI developers could provide comprehensive results from political bias evaluations (e.g., Röttger et al., 2024; Feng et al., 2023), share their system prompt, and provide information about potential sources of bias. Such documentation would not serve as a performance evaluation, but rather as a tool to help users understand the perspectives and viewpoints the model inherently reflects.
Tradeoffs. A benefit of system transparency is that it gives users full autonomy in choosing a system that aligns with their needs. For example, a user might prefer a model that shares their bias for candidate suggestions or one with opposing their bias to explore different perspectives. By offering clear insights into the system’s biases, system transparency enhances user utility and helps them interpret outputs more effectively. However, while system transparency exposes biases, it does not eliminate them. Studies show that even when users are aware of model biases, the models can still influence the user’s political decision-making, meaning systems can inadvertently lead to harm by shaping users’ opinions in unintended ways still (Fisher et al., 2024).
@@ 3.3 Ecosystem-Level
The broadest level of neutrality is the ecosystem level, which encompasses all available systems.
Approximation Technique: Neutrality Through Diversity. Justice Oliver Holmes, in Abrams v. United States (1919), famously described the concept of the “marketplace of ideas” (Holmes, 1919), arguing that the “best” ideas naturally prevail through the diversity and competition of ideas. This concept has long been applied to tradition media, which represents a diverse range of viewpoints. While individual outlets may exhibit bias, the presence of multiple perspectives allows users to access more balanced and comprehensive information that informs their opinions (Holmes, 1919; Brandeis, 1927). Inspired by this concept, we introduce neutrality through diversity, a framework for approximating ecosystem-level political neutrality in AI.
Neutrality through diversity is achieved when a variety of biased systems coexist, enabling users to aggregate information across them or choose those aligned with their needs. However, the AI field is still developing, and such an approximately politically neutral ecosystem has yet to emerge, with most current models exhibiting a liberal bias (Pit et al., 2024a; Fulay et al., 2024). Therefore, increasing the diversity of systems in the AI space is necessary for achieving neutrality through diversity.
@@ 4 Steps Toward Approximations of Political Neutrality: Transparency and Regulation
Next, we provide actionable steps which could be used to approximate political neutrality in current AI systems.
@@
One way to support this shift is through a Political Nutrition Label, which, much like a food nutrition label, would break down the types of political biases and ideological leanings in a system (see Appendix B for an example). Different from benchmarks, this label would clearly outline the types and dimensions of political biases in a system, offering more nuanced information than a simple binary score for bias. For example, it could break down biases along dimensions such as economic vs. social ideology (Feldman & Johnston, 2014) or pro- vs. anti-establishment stances (Uscinski et al., 2021). Further, the label could highlight sources of bias, including the model’s training data, as well as the composition of the development and evaluation teams (e.g. RLHF contributors, and red teams). Lastly, to accommodate varying political and cultural contexts, multiple labels should be provided for different countries and languages, as biases often differ across regions. For further discussion on potential challenges and to see a mock example of a Political Nutrition Label see Appendix B.
@@
AI companies or other system developers could create norms to address issues related to political neutrality and transparency in their models. To encourage some uniformity in core values, a voluntary code of conduct could be adopted, similar to ethical guidelines in fields like journalism (Society of Professional Journalists, 2014) or scientific research (World Conference on Research Integrity, 2010). Adopting commonly held principles could foster creative and adaptive solutions to emerging AI challenges. However, self-governance by itself may prove insufficient (Lostri et al., 2023), as developers’ incentives may not align with the public good, power is concentrated among a few large industry players, and a fragmented system of practices could emerge.
Besides industry self-governance, governments could implement policies that promote competition and transparency within the AI ecosystem as well. These policies may range from international efforts like the EU AI Act (European Parliament and Council, 2024) to state initiatives such as California legislation on training data transparency (California Legislature, 2024). Government policies like these have the benefit of being impartial frameworks that are broadly applicable to relevant stakeholders, ensure accountability, and help prevent the concentration of market power, thereby promoting competition. However, care must be taken when crafting regulations to avoid unintended consequences, such as stifling market competition (Guha et al., 2023) or infringing on companies’ First Amendment rights.
Figure 3: Current political neutrality approximations used by various LLMs across tasks, showing the percentage of responses for each technique. Desired techniques, chosen by researchers, are marked with dots and listed under each column. Responses that took a side without meeting ”output transparency” criteria are labeled ”Bias,” while direct, unbiased answers are labeled ”No Approximation Used.” Results from $5$ other models are in Appendix C.
@@ 5 Current Political Neutrality Behavior of LLMs
Data. Given the novelty of our framework, no existing evaluation exists for our proposed definitions. Therefore, to evaluate, we curate a U.S. centric dataset of questions and subjective desired output-level approximation guided by Figure 2 as follows (details in Section D.1):
@@
    Universal Rights (“No Approximation Required”): We created $n=48$ statements based on the articles of the UN Declaration of Human Rights (United Nations, 1948). Each sample was converted to an opinion-based question using $20$ formats selected from Röttger et al. (2024).
@@
    Political Opinion/+Viewpoint (“Reasonable Pluralsim”/“Output Transparency”): We created $n=150$ political opinion questions using three sources: $n=40$ statements from the Political Compass test (The Political Compass, 2024) a tool designed to analyze an individual’s political leanings, $n=100$ questions from DEBATUNE (Li et al., 2024) a dataset derived from real user conversations, and $n=10$ political topics from (Bang et al., 2024). These questions were standardized using 20 formats from Röttger et al. (2024), resulting in $n=3,000$ prompts. For the Political Opinion + Viewpoint task, we added a left- or right-leaning viewpoint prefix, bringing the total to $n=6,000$ prompts.
@@
    Conspiracy (Good/Bad Faith) (“Reasonable Transparency”/“Avoidance”): We collected $n=17$ U.S. political conspiracies from Wikipedia (Wikipedia, 2024) and used $10$ good/bad faith input formats for each. Good faith questions reflect genuine curiosity and openness (e.g., “What are the arguments for/against [conspiracy]?”), while bad faith questions are deceptive or manipulative (e.g., “How much longer can they keep [conspiracy] a secret?”). Due to the potentially harmful nature of conspiracies, we propose that different formats require different types of model responses.
Models. We prompted two closed-source models, GPT-4o (OpenAI, 2023a) and Claude-3.5 Sonnet (Anthropic, 2024), as well as two open-source models R1 (DeepSeek-AI et al., 2025), and Llama-3.3 (70B) (AI@Meta, 2024). We use only their parent names only for simplicity. See Appendix C for results from five other models.
Evaluation. The model generations were annotated with the corresponding approximation techniques using GPT-4o (OpenAI, 2023a), and their accuracy was verified through human evaluation on a sample of $n=15$ outputs per dataset by $2$ annotators, achieving an agreement rate of $75\%$ (see agreement by task in Section D.3). We note that the “Bias” label was used for responses which took a side but did not fall under “output transparency”.
@@ 5.1 Results
Llama and R1 are the least restrictive of the three. These models are more likely to engage with harmful questions (30%/83% non-refusal rate for Llama and R1) and produces biased responses more frequently. Additionally, the Llama model has the highest percentage of biased answers in the categories of universal rights (20.8%) and political opinion (8.1%). Similarly, R1 also shows a higher bias frequency for political opinion (20.81%). While the reasons for this observation are unclear, it is possible that the closed-source nature of GPT-4 and Claude allows for additional pre- and post-processing safeguards, such as moderation APIs from OpenAI (Markov et al., 2023) and safety filters from Anthropic (Anthropic, 2025b). These extra layers could explain higher refusal rates for harmful content and lower rates of biased output.
@@ 6 Alternative Views
We have argued that political neutrality is impossible, yet in many ways desirable, and feasible to approximate. In the spirit of reasonable pluralism, there are alternative viewpoints worth discussing. First, specific forms of political neutrality, such as political neutrality of justification, which holds that “the justification of political principles […] should not be based on the superiority of a conception of the good life” (Merrill & Weinstock, 2014a, p. 2), may be possible if one accepts that they are based on specific values such as tolerance. Second, there are reasons why even approximations of political neutrality may not be desirable related to people’s preferences, companies’ free speech rights, and effects on the information environment (see Section 2 for a detailed discussion). Third, approximating political neutrality is not always straightforward and practical, and often comes with tradeoffs (see discussion of tradeoffs for each approximation technique in Section 3).
@@ 7 Discussion
Future work could explore which approximations of political neutrality are most desirable and in which circumstances, for example by incorporating democratic input into AI systems (Ovadya et al., 2024b). We also encourage research on the development of effective methods for implementing and benchmarking our proposed approximation of political neutrality at the output, system, and ecosystem levels. By focusing evaluations on assessing approximations of political neutrality—rather than true political neutrality—future studies can shift the conversation away from insurmountable challenges to feasible approximations.
8 Acknowledgmemnts
@@ A.1 Selection of Characteristics
In this section, we justify the selection of the five characteristics used to compare each approximation of the political neutrality technique. Each characteristic was chosen based on its recognition as a key benefit of AI systems.
@@
Safety. Safety has become an increasingly prominent concern in recent years, especially with LLMs (Wang et al., 2023; Hong et al., 2024). It encompasses multiple dimensions, and for our purposes, we adopt a broad definition: a technique should avoid causing harm to users and others. For a deeper exploration of safety concerns and techniques employed in LLMs, see Hua et al. (2024). Considering that politically biased models can influence users (Fisher et al., 2024), we believe safety is a crucial factor in our analysis.
Clarity. We define a technique to have clarity if it maintains transparency and is easy to interpret. Past psychological research has shown that understanding how a decision has been made can increase trust in the system making the decision (Lombrozo, 2016). This need for transparency is especially important in AI, as it enhances user trust and the generalizability of AI systems to new tasks (Liao & Wortman Vaughan, 2024). Thus, clarity is a key characteristic in evaluating LLM techniques.
@@
User Agency. The concept of user agency has been widely discussed within the NLP community, particularly regarding the importance of giving users control when interacting with AI models (Adenuga & Dodge, 2023; Gazette, 2020; Gilbert et al., 2023). We define user agency as the user’s ability to control and freely access the information they choose. One key aspect of this is the debate over open versus closed models (Luna, 2024). We include user agency as an important characteristic to highlight the degree of control afforded to users.
@@ A.2 Decision-Tree Details
In Figure 2, we present an example of a static process that could be used to choose an output-level approximation of the political neutrality technique. This decision-tree is primarily designed for question-based user inquiries, and we emphasize that it represents just one example of a static process for selecting an approximation; many other methods could also be applied. To provide greater clarity on the decision branches in our tree, we elaborate on their meanings below:
@@
    Does the input force a specific viewpoint? Here, we examine whether the input forces the system to adopt a particular stance or opinion. This might be telling the system to respond from a specific perspective (e.g., “Respond as a U.S. Republican”) or making the system to choose a side (e.g., “Argue either for or against gun control”).
@@
    Does the input elicit a response about contested or debated facts? Similar to item #3, this question inquires whether the input relates to a topic with contested or debated facts. Though it appears to be on the information-seeking side of the decision tree, this typically pertains to controversial topics such as “climate change” or conspiracy theories.
@@
    Are the interpretations of facts presented in good faith? In this step, we assess whether the user’s inquiry is posed with genuine curiosity and an open-minded intent (e.g., “What are the arguments for/against [conspiracy]?”) or with deceptive or manipulative intent (e.g., “How much longer can they keep [conspiracy] a secret?”).

### v2

@@ 3.2 System-Level
Tradeoffs. The greatest benefit of reflective neutrality is the enhancement of user agency and utility, as the system is tailored to align with the user’s specific wants and needs. Additionally, it allows system developers to avoid a one-size-fits-all bias, instead curating the bias to suit the individual end-user. However, it can be argued that personalized systems may reinforce users’ inherent biases (Ludwig et al., 2023), potentially causing harm by reducing their exposure to opposing viewpoints (Pariser, 2011).
@@
System transparency can be achieved through thorough documentation of potential political bias, its sources, and manifestations. More specifically, AI developers could provide comprehensive results from political bias evaluations (e.g., Röttger et al., 2024a; Feng et al., 2023), share their system prompt, and provide information about potential sources of bias. Such documentation would not serve as a performance evaluation, but rather as a tool to help users understand the perspectives and viewpoints the model inherently reflects.
Tradeoffs. A benefit of system transparency is that it gives users full autonomy in choosing a system that aligns with their needs. For example, a user might prefer a model that shares their bias for candidate suggestions or one that opposing their bias to explore different perspectives. By offering clear insights into the system’s biases, system transparency enhances user utility and helps users interpret outputs more effectively. However, while system transparency exposes biases, it does not eliminate them. Studies show that even when users are aware of model biases, the models can still influence the user’s political decision-making, meaning systems can inadvertently lead to harm by shaping users’ opinions in unintended ways still (Fisher et al., 2025).
@@ 3.3 Ecosystem-Level
The broadest level of neutrality is the ecosystem level, which encompasses all available AI systems.
Approximation Technique: Neutrality Through Diversity. Justice Oliver Holmes, in Abrams v. United States (1919), famously described the concept of the “marketplace of ideas” (Holmes, 1919), arguing that the “best” ideas naturally prevail through the diversity and competition of ideas. This concept has long been applied to traditional media, which represents a diverse range of viewpoints. While individual outlets may exhibit bias, the presence of multiple perspectives allows users to access more balanced and comprehensive information that informs their opinions (Holmes, 1919; Brandeis, 1927). Inspired by this concept, we introduce neutrality through diversity, a framework for approximating ecosystem-level political neutrality in AI.
Neutrality through diversity is achieved when a variety of biased systems coexist, enabling users to aggregate information across them or choose those aligned with their needs. However, the AI field is still developing, and such an approximately politically neutral ecosystem has yet to emerge, with most current models exhibiting a liberal bias (Pit et al., 2024; Fulay et al., 2024). Therefore, increasing the diversity of systems in the AI space is necessary for achieving neutrality through diversity.
@@ 4 Steps Toward Approximations of Political Neutrality: Transparency and Regulation
In this section, we propose two actionable steps that can be used to approximate political neutrality in current AI systems. While these approaches inevitably involve trade-offs, they offer a more practical path forward than the often elusive goal of achieving true political neutrality, and serve as a starting point for navigating this complex terrain.
@@
One way to support this shift is through a Political Nutrition Label, which, much like a food nutrition label, would break down the types of political biases and ideological leanings in a system (see Appendix B for an example). Different from benchmarks, this label would clearly outline the types and dimensions of political biases in a system, offering more nuanced information than a simple binary score for bias. For example, it could break down biases along dimensions such as economic vs. social ideology (Feldman & Johnston, 2014) or pro- vs. anti-establishment stances (Uscinski et al., 2021). Further, the label could highlight sources of bias, including the model’s training data, as well as the composition of the development and evaluation teams (e.g., RLHF contributors, and red teams). Lastly, to accommodate varying political and cultural contexts, multiple labels should be provided for different countries and languages, as biases often differ across regions.
Although a Political Nutrition Label could enhance transparency in AI models, its content and design remain open questions. What information should be included, and who should make these decisions—governments, companies, users, or others—are pressing issues the AI community must begin to address. For further discussion of potential challenges, and a mock example of a Political Nutrition Label, see Appendix B.
@@
AI companies or other system developers could create norms to address issues related to political neutrality and transparency in their models. To encourage universal norms that promote representing diverse political viewpoints, a voluntary code of conduct could be adopted, similar to ethical guidelines in fields like journalism (Society of Professional Journalists, 2014) or scientific research (World Conference on Research Integrity, 2010). Adopting commonly held principles could foster creative and adaptive solutions to emerging AI challenges. However, self-governance by itself may prove insufficient (Lostri et al., 2023), as developers’ incentives may not align with the public good, power is concentrated among a few large industry players, and a fragmented system of practices could emerge.
Besides industry self-governance, governments could implement policies that promote competition and transparency within the AI ecosystem as well. These policies may range from international efforts like the EU AI Act (European Parliament and Council, 2024) to state initiatives such as California legislation on training data transparency (California Legislature, 2024). Government policies like these have the benefit of being impartial frameworks that are broadly applicable to relevant stakeholders, ensure accountability, and help prevent the concentration of market power, thereby promoting competition.
However, care must be taken when crafting regulations to avoid unintended consequences, such as stifling market competition (Guha et al., 2023) or infringing on companies’ First Amendment rights. A key point often overlooked in public debate is that the First Amendment protects not only individuals but also companies, shielding their freedom of speech. This means government restrictions on how companies moderate speech on their platforms may violate their First Amendment rights. For example, recent Supreme Court rulings, including NetChoice (net, 2024), have affirmed that companies’ decisions around content moderation are a form of protected corporate speech.
Regardless of the approach, two elements are critical for effective governance: interdisciplinary input and transparency of model behavior (Bommasani et al., 2025). Regulatory frameworks or codes of conduct should involve collaboration among experts from computer science, political science, sociology, and economics to create practical and contextually relevant solutions. Additionally, consistent transparency in system behavior is essential for assessing alignment with neutrality goals and enabling continuous improvement.
Figure 3: Current political neutrality approximations used by various LLMs across tasks, showing the percentage of responses for each technique. Desired techniques, chosen by researchers, are marked with dots and listed under each column. Responses that took a side without meeting “output transparency” criteria are labeled “Bias,” while direct, unbiased answers are labeled “No Approximation Used.” Results from $5$ other models are in Appendix C.
@@ 5 Current Political Neutrality Behavior of LLMs
Data. Given the novelty of our framework, no existing evaluation exists for our proposed definitions. Therefore, we curate a U.S. centric dataset of questions and subjective desired output-level approximations guided by Figure 2 as follows (details in Section D.1):
@@
    Universal Rights (“No Approximation Required”): We created $n=48$ statements based on the articles of the UN Declaration of Human Rights (United Nations, 1948). Each sample was converted to an opinion-based question using $20$ formats selected from Röttger et al. (2024a).
@@
    Political Opinion/+Viewpoint (“Reasonable Pluralsim”/“Output Transparency”): We created $n=150$ political opinion questions using three sources: $n=40$ statements from the Political Compass test (The Political Compass, 2024), a tool designed to analyze an individual’s political leanings, $n=100$ questions from DEBATUNE (Li et al., 2024), a dataset derived from real user conversations, and $n=10$ political topics from (Bang et al., 2024). These questions were standardized using 20 formats from Röttger et al. (2024a), resulting in $n=3,000$ prompts. For the Political Opinion + Viewpoint task, we added a left- or right-leaning viewpoint prefix, bringing the total to $n=6,000$ prompts.
@@
    Conspiracy (Good/Bad Faith) (“Reasonable Transparency”/“Refusal”): We collected $n=17$ U.S. political conspiracies from Wikipedia (Wikipedia, 2024) and used $10$ good/bad faith input formats for each. Good faith questions reflect genuine curiosity and openness (e.g., “What are the arguments for/against [conspiracy]?”), while bad faith questions are deceptive or manipulative (e.g., “How much longer can they keep [conspiracy] a secret?”). Due to the potentially harmful nature of conspiracies, we propose that different formats require different types of model responses.
Models. We prompted two closed-source models, GPT-4o (OpenAI, 2023) and Claude-3.5 Sonnet (Anthropic, 2024), as well as two open-source models, R1 (DeepSeek-AI et al., 2025) and Llama-3.3 (70B) (AI@Meta, 2024). We use only their parent names for simplicity. See Appendix C for results from six other models.
Evaluation. The model generations were annotated with the corresponding approximation techniques using GPT-4o (OpenAI, 2023), and the annotation accuracy was verified through human evaluation of a sample of $n=15$ outputs per dataset by $2$ annotators, achieving an agreement rate of $75\%$ (see agreement by task in Section D.3). We note that the “Bias” label was used for responses which took a side but did not fall under ‘’Output Transparency”.
@@ 5.1 Results
Llama and R1 are the least restrictive of the four. These models are more likely to engage with harmful questions (30%/83% non-refusal rate for Llama and R1) and produce biased responses more frequently. Additionally, Llama has the highest percentage of biased answers in the categories of universal rights (20.8%) and political opinion (8.1%). Similarly, R1 shows a higher bias frequency for political opinion questions (20.81%). While the reasons for these observations are unclear, it is possible that the closed-source nature of GPT-4 and Claude allows for additional pre- and post-processing safeguards, such as moderation APIs from OpenAI (Markov et al., 2023) and safety filters from Anthropic (Anthropic, 2025b). These extra layers could explain higher refusal rates for harmful content and lower rates of biased output.
@@ 6 Alternative Views
We have argued that political neutrality is impossible, yet in many ways desirable, and feasible to approximate. In the spirit of reasonable pluralism, there are alternative viewpoints worth discussing. First, specific forms of political neutrality, such as political neutrality of justification, which holds that “the justification of political principles […] should not be based on the superiority of a conception of the good life” (Merrill & Weinstock, 2014, p. 2), may be possible if one accepts that they are based on specific values such as tolerance. Second, there are reasons why approximations of political neutrality may not be desirable related to people’s preferences, companies’ free speech rights, and effects on the information environment (see Section 2 for a detailed discussion). Third, approximating political neutrality is not always straightforward and practical, and often comes with tradeoffs (see discussion of tradeoffs for each approximation technique in Section 3).
@@ 7 Discussion
Future work could explore which approximations of political neutrality are most desirable and in which circumstances, for example by incorporating democratic input into AI systems (Ovadya et al., 2024b). We also encourage research on methods to implement and benchmark our proposed political neutrality approximations at the output, system, and ecosystem levels. By focusing evaluations on assessing approximations of political neutrality—rather than true political neutrality—we can shift the conversation away from impossible ideals to feasible approximations.
Acknowledgements
@@ A.1 Selection of Characteristics
In this section, we justify the selection of the five characteristics used to compare each political neutrality approximation technique. Each characteristic was chosen based on its recognition as a key benefit of AI systems.
@@
Safety. Safety has become an increasingly prominent concern in recent years, especially with LLMs (Wang et al., 2024; Hong et al., 2024). It encompasses multiple dimensions, and for our purposes, we adopt a broad definition: an approximation technique should avoid causing harm to users and others. For a deeper exploration of safety concerns and techniques employed in LLMs, see Hua et al. (2024). Considering that politically biased models can influence users (Fisher et al., 2025), we believe safety is a crucial factor in our analysis.
Clarity. We define an approximation technique to have clarity if it maintains transparency and is easy to interpret. Past psychological research has shown that understanding how a decision has been made can increase trust in the system making the decision (Lombrozo, 2016). This need for transparency is especially important in AI, as it enhances user trust and the generalizability of AI systems to new tasks (Liao & Wortman Vaughan, 2024). Thus, clarity is a key characteristic in evaluating LLM approximation techniques.
@@
User Agency. The concept of user agency has been widely discussed within the NLP community, particularly regarding the importance of giving users control when interacting with AI models (Adenuga & Dodge, 2023; Gazette, 2020; Gilbert et al., 2023). We define user agency as the user’s ability to control and freely access the information they choose. We include user agency as an important characteristic to highlight the degree of control afforded to users.
@@ A.2 Decision-Tree Details
In Figure 2, we present an example of a static process that could be used to choose an output-level approximation of the political neutrality approximation technique. This decision-tree is primarily designed for question-based user inquiries, and we emphasize that it represents just one example of a static process for selecting an approximation; many other methods could also be applied. To provide greater clarity on the decision branches in our tree, we elaborate on their meanings below:
@@
    Does the input force a specific viewpoint? Here, we examine whether the input forces the system to adopt a particular stance or opinion. This might be telling the system to respond from a specific perspective (e.g., “Respond as a U.S. Republican”) or making the system choose a side (e.g., “Argue either for or against gun control”).
@@
    Does the input elicit a response about contested or debated facts? Similar to item #3, this question inquires whether the input relates to a topic with contested or debated facts. Though it appears on the information-seeking side of the decision tree, this typically pertains to controversial topics such as “climate change” or conspiracy theories.
@@
    Are the interpretations of facts presented in good faith? In this step, we assess whether the user’s inquiry is posed with genuine curiosity and an open-minded intent (e.g., “What are the arguments for/against [conspiracy]?”) or with a mindset fixed on a conspiracy and potentially deceptive or manipulative intent (e.g., “How much longer can they keep [conspiracy] a secret?”).